# SMILES has to go : Representation of Molecules via Algebraic Data Types

### v1

@@ 3.3.1 Example of Orbitals : Hydrogen, Carbon & Oxygen
3.4 Reaction Networks, Group Theory and Serialisation
3.4.1 Reaction Networks
@@
Here’s an example of how the Reaction data type can be defined:
data Reaction = Reaction {
    reactants :: [(Double, Molecule)]
  , products :: [(Double, Molecule)]
  , conditions :: [Condition]
  , reactionRate :: Double
data Condition = TempCondition {temperature :: Double, times :: Times}
               | PressureCondition {pressure :: Double, times :: Times}
newtype Times = Times (StartTime Double, EndTime Double)
@@
newtype ChemicalReactionNetwork = ChemicalReactionNetwork {
    reactions :: [Reaction]
@@
exampleReaction :: Reaction
exampleReaction = Reaction {
    reactants = [(2.0, h2), (1.0, o2)]
  , products = [(2.0, h2o)]
  , conditions =
    [ TempCondition 500.0 (Times (0.0, 10.0))
    , PressureCondition 1.0 (Times (0.0, 10.0))
    reactionRate = 0.1
@@
3.4.2 Group Theory
Moreover, the ADT can be instantiated as a group to incorporate group theory concepts, which play a crucial role in understanding molecular symmetry and its impact on chemical properties. In particular, molecules can be made instances of rotation groups, such as the cyclic group $C3$, which represents the rotational symmetry of molecules with threefold rotational axes.
Here’s an example of how a MoleculeRotation data type can be defined and made an instance of the Group typeclass:
class Group g where
    mul :: g -> g -> g     -- Group operation: combining two operations
    inv :: g -> g          -- Inverse of an operation
    e   :: g -> g          -- Identity operation
data MolRot = MolRot Molecule Coordinate Double deriving (Show)
instance Group MolRot where
    mul (MolRot mol1 axis1 angle1) (MoleculeRotation mol2 axis2 angle2) =
        combineRotations mol1 axis1 angle1 axis2 angle2
    inv (MoleculeRotation mol axis angle) = MolRot mol axis (-angle)
    e (MoleculeRotation mol axis angle) = MolRot mol (Coordinate 0 0 0) 0
In this representation, a MolRot consists of a Molecule, a rotation axis specified by a Coordinate, and a rotation angle (in radians). The Group typeclass defines the necessary operations for a group, namely the group operation (mul), the inverse operation (inv), and the identity operation (e).
@@ 3.4.2 Group Theory
3.4.3 Reading & Writing
{verbnobox}
[] – Writing the molecule to a file writeMoleculeToFile :: FilePath -> Molecule -> IO () writeMoleculeToFile filePath molecule = writeFile filePath (show molecule)
– Reading the molecule from a file readMoleculeFromFile :: FilePath -> IO Molecule readMoleculeFromFile filePath = do contents <- readFile filePath return (read contents)
– Example usage main :: IO () main = do – Write the methane molecule to a file writeMoleculeToFile "methane.hs" methane
– Read the methane molecule from the file molecule <- readMoleculeFromFile "methane.hs"
– Print the molecule read from the file print molecule
Figure 9: The provided Haskell code demonstrates how molecules can be easily read from and written to files.
The code defines two functions:
  1. 1.
    writeMoleculeToFile :: FilePath -> Molecule -> IO () – This function takes a file path and a Molecule as input and writes the molecule to the specified file. It uses the show function to convert the molecule to a string representation and then writes it to the file using the writeFile function.
  2. 2.
    readMoleculeFromFile :: FilePath -> IO Molecule – This function takes a file path as input and reads a Molecule from the specified file. It uses the readFile function to read the contents of the file as a string and then uses the read function to parse the string back into a Molecule.
The main function provides an example usage of these functions. It first writes the methane molecule to a file named methane.hs using the writeMoleculeToFile function. Then, it reads the molecule back from the same file using the readMoleculeFromFile function and prints the molecule to the console.
@@ 3.4.3 Reading & Writing
This functionality is particularly useful when working with large molecules or when you need to persist molecules across different program runs. It provides a convenient way to save and load molecular data, enabling further analysis, visualization, or processing of the molecules as needed.
3.4.4 A simple example of probablistic programming
Consider the following code:
{verbnobox}
[] – Observed molecule (e.g., H2O) observedMolecule :: Molecule observedMolecule = h2o
– A generative model for a molecule moleculeModel :: Meas Molecule moleculeModel = do let numAtoms = 3 atoms <- replicateM numAtoms \$ do symbol <- sample \$ uniformD [C, N, O, H] coord <- sampleCoordinate let attr = elementAttributes symbol let shellsVar = elementShells symbol return \$ Atom atomID = 0, atomicAttr = attr, coordinate = coord, shells = shellsVar let molecule = Molecule atoms = atoms, bonds = randomBonds 3
– Score the model based on the distance to the observed molecule let distance = distMolecule molecule observedMolecule score \$ normalPdf 0 1 distance
return molecule
– Sample a coordinate from a normal distribution sampleCoordinate :: Meas Coordinate sampleCoordinate = do x <- sample \$ normal 0 1 y <- sample \$ normal 0 1 z <- sample \$ normal 0 1 return \$ Coordinate x y z
– Calculate the Euclidean distance between two coordinates euclideanDistance :: Coordinate -> Coordinate -> Double euclideanDistance (Coordinate x1 y1 z1) (Coordinate x2 y2 z2) = sqrt \$ (x1 - x2)^2 + (y1 - y2)^2 + (z1 - z2)^2
– Calculate the distance between two molecules distMolecule :: Molecule -> Molecule -> Double distMolecule mol1 mol2 = let coords1 = map coordinate (atoms mol1) coords2 = map coordinate (atoms mol2) distances = [euclideanDistance c1 c2 | c1 <- coords1, c2 <- coords2] in minimum distances
– Example usage main :: IO () main = do samples <- mh 0.1 moleculeModel let (molecules, weights) = unzip \$ take 1000 \$ drop 1000 samples putStrLn \$ "Sampled molecules: " ++ show molecules putStrLn \$ "Weights: " ++ show weights
Figure 10: Haskell code for generating and conditioning molecules using a probabilistic model. The model generates molecules and scores them based on their distance to an observed molecule (e.g., H2O). Metropolis-Hastings sampling is used to generate samples from the conditioned model.
Explanation:
@@ 4 Discussion
In a recent review of molecular representation, four main types of molecular representation are considered [Wigh et al., 2022]. These are Strings, Graphs, Fingerprints and Variational Auto Encoders. Here, relative advantages and disadvantages of all four approaches will be investigated and evaluated against the proposal here.
4.0.1 Unsorted Advantages compared to other representaions
In the Molecule ADT, whilst the same problem may exist with atom IDs for instance, there is a structured way to represent symmetries with group theoretic typeclasses.
4.1 The three types of representation
4.1.1 Storage
The ADT representation may have some drawbacks in terms of storage efficiency compared to more compact file formats like molfiles. The verbose nature of the ADT syntax could result in larger file sizes, especially for large molecular datasets. This could impact storage requirements and data transfer times. Nevertheless, the benefits of the ADT’s expressiveness and extensibility may outweigh the storage efficiency concerns in many cases, particularly when the accurate representation of complex molecular features is crucial.
4.1.2 Transmission
For transmission purposes, the ADT representation can be serialized into a human-readable format using the show and read functions in Haskell. This allows for the easy exchange of molecular data between cheminformaticians using plain text formats, such as .hs files. In the context of molecular representation transmission, it is important to distinguish between the serialized format of the data and the underlying data type itself. While the proposed algebraic data type (ADT) representation can be serialized into a human-readable string format using the show and read functions in Haskell, it is crucial to recognize that the ADT itself is not merely a string.
The ADT representation is a structured and type-safe data type that encapsulates the complex features and properties of molecules. The type annotations within the ADT provide a rich semantic context that goes beyond the limitations of plain strings. These type annotations ensure that the data is well-formed, consistent, and adheres to the specified structure and constraints of the molecular representation.
In the modern era of cheminformatics, it is essential to distinguish between the data itself and the type of data being used. Treating molecular representations as mere strings can lead to ambiguity, inconsistency, and potential errors in interpretation and manipulation. Strings lack the inherent structure and semantics that are crucial for accurately capturing and reasoning about molecular properties and behaviors.
By leveraging the power of ADTs and type systems, cheminformaticians can work with molecular representations in a more precise, reliable, and meaningful way. The type annotations serve as a form of documentation and validation, ensuring that the data conforms to the expected structure and constraints. This enables more robust and error-free manipulation of molecular data, reducing the chances of introducing inconsistencies or invalid states.
Moreover, the use of ADTs facilitates the development of type-safe and composable functions and algorithms that operate on molecular data. Type-driven development allows for the creation of reusable and modular code, promoting code clarity, maintainability, and scalability. This is particularly valuable in the context of complex cheminformatics workflows, where the ability to reason about the types and structure of data can lead to more efficient and reliable implementations.
In contrast, treating molecular representations as plain strings can hinder the development of robust and maintainable cheminformatics software. Strings provide no guarantees about the structure or validity of the data, leading to the need for extensive runtime checks and error handling. This can result in more brittle and error-prone code, making it harder to reason about the correctness and behavior of the software. Therefore, in this age of advanced computational tools and techniques, it is imperative to embrace the distinction between data and the type of data being used.
4.1.3 Computational
The ADT representation offers several computational advantages over other molecular representations like strings, graphs, bond electron matrices, and fingerprints. The structured nature of the ADT allows for efficient and type-safe manipulation of molecular data within a programming environment. The use of algebraic data types enables pattern matching and recursive processing, which can lead to concise and expressive code for complex molecular operations. The strong static typing of the ADT representation catches potential errors at compile-time, reducing the chances of runtime failures and increasing the overall reliability of cheminformatics software. In contrast, string-based representations like SMILES or SELFIES may require additional parsing and validation steps to ensure the correctness of the molecular structure. Graph-based representations, while capturing the connectivity of atoms, may lack the expressiveness to represent more nuanced features like stereochemistry or resonance structures. Bond electron matrices and fingerprints, although useful for specific applications like similarity searching or machine learning, may not provide a complete and detailed representation of the molecular structure.
Apart from the obvious benefits of being able to represent molecules which cannot currently be represented using string based representations (even with the creation of new and complicated grammars), our ADT confers other crucial benefits:
  1. 1.
    By directly synthesizing molecules in accordance with the grammar defined by the ADT, machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation into another format.
  2. 2.
    Our ADT is relatively extensible, meaning that the cheminformatics community can modify the grammar of the ADT in accordance with their needs, (e.g. to represent polymers or tautomeric molecules).
4.2 Strings
4.2.1 The future of SELFIES and of molecular representation
Some of the authors of SELFIES happen to have produced [Krenn et al., 2022], which points to future representational projects for molecules. In the future of SELFIES, [Krenn et al., 2022] they suggest that strings can represent Turing complete programming languages. Why not just drop the word strings (which have limitations as a data type) and use pre-existing well developed programming languages? They title project 8 as; “Future Project 8: Introducing “molecular programming languages” which they mention means adding 3D information to the representation, with a meta ‘for’ keyword, so despite the title, is in no way an actual suggestion of a practical molecular programming language, in the usual sense of programming language.
The authors of [Krenn et al., 2022] describe programming languages as strings, but indeed more formally they are sets of strings. Clarification may be worthwhile to ensure integrity of discourse between communities. Strings in this article are a data type that support a very limited number of operations, such as [CharAt, Concatenation, Contains, Equality, Find, indexof,
instrlength, locate, Lowercase, partition, replace, reverse, split]. Whilst software engineers write programs with text, it would be misleading to suggest that one reflects on applying string operations to the text of our programs. This would likely break the syntax of the code, rendering the code without formal denotation, let alone well-formed structure. The programmer thinks about and reasons on the object level of the program itself in accordance with some framework determined by the Church-Turing thesis, not on its textual syntactic representation. They mention “Python is of course a Turing-complete language, which means, strings can encode the most powerful computational algorithms.” which suggests strings are of prime importance [^7] The authors also write like that in this substack article: https://aspuru.substack.com/p/molecular-graph-representations-and. Looking at programming languages through the Church Turing thesis, or data and functions which reflects the division that Von Neumann suggested, is more faithful to the jargon used by the programming languages community, and can help to avoid confusion. It is important to remember that all of the “power” of strings occurs with Abacus Machines or the lambda calculus, which is not just a string. Not to mention that one dimensional strings are difficult to read. They also suggest strings as the “esperanto” of chemistry, which seems absurd thinking about strings such as those in Figure LABEL:morphine.
They also propose “BEYOND ORGANIC CHEMISTRY: COMPLICATED BONDS”, which advocates using the Dietz representation, which is precisely the one in use in this articles ADT.
They also suggest “Future Project 9: A 100% robust programming language” by finding a syntax for a programming language that however you mutate the source program code, a valid program is always produced. It seems unclear how to do this except with a trivial interpreter or parser. It seems this is an attempt to generalise their own robustness of mutation to the most general data structure; programs themselves. Despite their title, this does not correspond to what has been proposed in this article.
5 Future Work
The representation proposed is for cheminformaticians, not chemists, and as such the representations used by chemists, need not even be viewed on computer screens. For chemists, inspiration can be taken from work on Tangible Bits or Morphological Computation [Ishii and Ullmer, 1997], resulting in virtual reality or high frequency sound waves to levitate balls and sticks representing Lewis structures [Marzo et al., 2015].
Addition of optional tautomeric data to the representation would be very helpful. In [Pospisil et al., 2003] and [Sitzmann et al., 2010] the importance of tautomers is made clear. In [Sitzmann et al., 2010], a database of 680 million tautomers is established. They are in medications from sildenafil (Viagra) [Pajka et al., 2022], to Warfarin, to Remdesivir (used for Covid 19), Tetracyclines and antibiotics. A central problem in drug design is predicting ligand binding to proteins [Boyles et al., 2020]. Tautomers are important because tautomers affect the conformation, which affects the protein ligand attachment. [Pospisil et al., 2003] states “Tautomers are often disregarded in computer-aided molecular modeling applications. Little is known about the different tautomeric states of a molecule and they are rarely registered in chemical databases. Tautomeric forms of a molecule differ in shape, functional groups, surface, and hydrogen-bonding pattern. Calculation of physical–chemical properties and molecular descriptors differ from one tautomeric state to the other as it is demonstrated with an example of the log P calculation, similarity index, and the complementarity pattern to the targeted protein. Considering tautomery in ligand–protein interactions therefore has a significant impact on the prediction of the ligand binding using various docking techniques. This article points [to] hitherto unaddressed issue of tautomerism in computer-aided drug design
One avenue for future work is the integration of the proposed algebraic data type (ADT) representation with geometric deep learning techniques to exploit the inherent symmetries and invariances present in molecular structures [Atz et al., 2021]. By leveraging the mathematical properties of ADTs, such as group invariances and equivariances and their verification in Haskell through typeclasses, one could develop neural network architectures that are specifically tailored to the unique characteristics of those molecules. For instance, the atomIDs used in the ADT representation are essentially arbitrary labels and should not influence the learning process. Neural networks can be designed to be invariant to permutations of atomIDs, ensuring that molecularly equivalent structures are treated identically regardless of the specific ID assignment. Similarly, certain molecular features, such as the atomSymbol, are both translation and rotation invariant, while others, may only be translation invariant. By incorporating these symmetries directly into the neural network architecture, we can significantly reduce the amount of training data required and improve the generalization capability of the models. Geometric deep learning techniques, such as equivariant neural networks and graph neural networks, have shown great promise in capturing the complex spatial and relational structure of molecules.
To further demonstrate the practical utility and superiority of the proposed algebraic data type (ADT) representation, an important direction for future work is to empirically validate its performance on real-world machine learning tasks in cheminformatics. By benchmarking the ADT representation against existing approaches, such as SMILES and SELFIES, on tasks like molecular property prediction, virtual screening, and de novo drug design, we can quantitatively assess its effectiveness and identify areas for further improvement. Moreover, the current ADT representation can be extended to incorporate additional quantum mechanical properties, such as spin and angular momentum, into the electronic configuration of atoms. This enhancement would enable a more accurate and comprehensive description of the electronic structure, potentially leading to improved predictions of molecular properties and reactivity. To further strengthen the type safety and invariance guarantees of the representation, exploring the use of dependently typed programming languages, such as Agda, could be highly beneficial. Dependent types allow for the encoding of more sophisticated invariants and constraints directly into the type system, ensuring that only valid and consistent molecular structures are expressible. This increased level of type safety can help catch potential errors at compile-time and provide stronger guarantees about the correctness of the representation. Another important extension to the ADT representation is the inclusion of polymers, which are crucial in many areas of chemistry and materials science. By leveraging the power of lists and recursive data structures, the representation can be augmented to handle polymeric molecules seamlessly. This extension would enable the modeling and analysis of complex macromolecular systems, opening up new possibilities for polymer design and optimization. Finally, to enhance the modularity and composability of the ADT representation, exploring techniques like modular syntax trees and the “Data Types a la Carte” approach [Swierstra, 2008] could prove highly valuable. By decomposing the molecular representation into smaller, reusable components, we can create a more flexible and adaptable framework. This modular design would allow for the easy integration of new features, the creation of domain-specific languages for molecular manipulation, and the development of reusable libraries and tools for cheminformatics. By pursuing these avenues of future work, we can further strengthen the capabilities and impact of the ADT-based molecular representation, positioning it as a powerful and versatile tool for advancing the field of cheminformatics and accelerating the discovery of novel molecules with desired properties.
Some of the most important drugs in history, such as Warfarin, Allopurinol (a medication to treat gout which has nine tautomers), and antibiotics, display tautomerism [Porter, 2010], but existing transmission representations such as SMILES, and others which rely on molecular graphs, cannot represent tautomeric molecules in a single structure. Furthermore graph structures alone cannot straightforwardly be extended to resonant structures, tautomers, 3D information and organometallics [David et al., 2020] & [Wigh et al., 2022]. Dietz [1995]. This could be useful further work.
Further work could explore several avenues to enhance the proposed algebraic data type (ADT) representation and its applications. Integrating Haskell’s Lens library [Steckermeier, 2015] could simplify the manipulation of complex molecular structures by providing a composable and modular way to access and modify nested data fields. Implementing backpropagation using differentiable programming [Elliott, 2009] would enable the training of deep learning models directly on the ADT representation, leveraging the structure and semantics of the molecular data. To ensure the practical value and completeness of the ADT, empirical verification through extensive feedback and collaboration with chemists and cheminformaticians is crucial. This process would involve presenting the ADT to domain experts, gathering their insights, and iteratively refining the representation based on their requirements. Assessing the ADT’s performance on a prediction task using the QM9 chemical benchmark dataset [Pinheiro et al., 2020] would provide valuable insights into its effectiveness in capturing relevant chemical information compared to existing state-of-the-art methods. Furthermore, integrating the ADT with quantum sampling techniques could lead to breakthroughs in predicting molecular properties by efficiently simulating complex quantum systems. However, realizing this goal would require close collaboration between experts in quantum computing, chemical informatics, and molecular modeling to develop efficient quantum circuits and optimize the mapping between the ADT and quantum hardware. These future directions highlight the potential of the ADT representation to revolutionize various aspects of computational chemistry and materials science.

### v2

@@ 46 Orbital Pz 1 (Just (Coordinate 0 0 1))])
6 Experiments and extensions
6.1 Reaction Networks
@@
Here’s an example of how the Reaction data type can be defined: 1data Reaction = Reaction { 2 reactants :: [(Double, Molecule)] 3 , products :: [(Double, Molecule)] 4 , conditions :: [Condition] 5 , reactionRate :: Double 6} 7 8data Condition = TempCondition {temperature :: Double, times :: Times} 9 | PressureCondition {pressure :: Double, times :: Times} 10 11newtype Times = Times (StartTime Double, EndTime Double) Listing 9: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate.
@@
1newtype ChemicalReactionNetwork = ChemicalReactionNetwork {
2 reactions :: [Reaction]
3}
Listing 10: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate.
@@
1exampleReaction :: Reaction
2exampleReaction = Reaction {
3 reactants = [(2.0, h2), (1.0, o2)],
4 products = [(2.0, h2o)],
5 conditions = [ TempCondition 500.0 (Times (0.0, 10.0)),
6 PressureCondition 1.0 (Times (0.0, 10.0)) ],
7 reactionRate = 0.1
8}
Listing 11: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate.
@@
6.2 Group Theory
Moreover, the ADT can be instantiated as a group to incorporate group theory concepts, which play a crucial role in understanding molecular symmetry and its impact on chemical properties. Exploiting these group symmetries can aid neural networks by the use of group equivariance and group invariance [2]. Invariances can be used also with atomic IDs meaning that the same atom represented with different ID’s can be quotiented and so treated as the same, allowing less data to be needed in order to train a neural network, for instance. In particular, molecules can be made instances of rotation groups, such as the cyclic group $C3$, which represents the rotational symmetry of molecules with threefold rotational axes.
Here’s an example of how a MoleculeRotation data type can be defined and made an instance of the Group typeclass: 1class Group g where 2 mul :: g -> g -> g -- Group operation: combining two operations 3 inv :: g -> g -- Inverse of an operation 4 e :: g -> g -- Identity operation 5 6data MolRot = MolRot Molecule Coordinate Double deriving (Show) 7 8instance Group MolRot where 9 mul (MolRot mol1 axis1 angle1) (MoleculeRotation mol2 axis2 angle2) = 10 combineRotations mol1 axis1 angle1 axis2 angle2 11 12 inv (MoleculeRotation mol axis angle) = MolRot mol axis (-angle) 13 14 e (MoleculeRotation mol axis angle) = MolRot mol (Coordinate 0 0 0) 0 Listing 12: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate. In this representation, a MolRot consists of a Molecule, a rotation axis specified by a Coordinate, and a rotation angle (in radians). The Group typeclass defines the necessary operations for a group, namely the group operation (mul), the inverse operation (inv), and the identity operation (e).
@@ 6.2 Group Theory
6.3 Storage, Transmission, and Readability
As well the computational purposes discussed elsewhere in the paper, the ADT provides advantages with respect to storage of the molecule, and transmission between cheminformaticians.
Molecules in the ADT can be serialised and stored as .hs files, which may be used for storage, and for transmission of molecular data, perhaps via large chemical databases, through straightforward use of the Show and Read instances in Haskell, as shown in Listing 14.
Additionally, as .hs files will be interpreted with Haskell, molecules can be straightforwardly compressed as Listing 13 demonstrates by using foldr to compress the bondsData list of the benzene molecule. Although this may raise the issue of “canonicalisation” for storage of molecules, it should be noted the bondsData in Listing 13 and Listing 6 will be stored the same way inside of a Haskell program, as the corresponding data is the same in each case. This could be beneficial in the representation of large molecules, as only a small number of bonds or atoms need be explicitly described to encode the molecule for storage.
1
2bondsData :: [((Int, Int), Bond)]
[…]
6 where
7 single = Bond {delocNum = 2, atomIDs = Nothing}
8 aromatic = Bond {delocNum = 6, atomIDs = Just [1,2,3,4,5,6]}
Listing 13: Using foldr to compress the bondsData of the molecule benzene in the ADT representation. As Benzene is symmetric, the bonding system can be specified by only describing three bonding systems, and using those to construct the bondsData list.
1-- Writing the molecule to a file
2writeMoleculeToFile :: FilePath -> Molecule -> IO ()
3writeMoleculeToFile filePath molecule = writeFile filePath (show molecule)
4
5-- Reading the molecule from a file
6readMoleculeFromFile :: FilePath -> IO Molecule
7readMoleculeFromFile filePath = do
8 contents <- readFile filePath
9 return (read contents)
10
11-- Example usage
12main :: IO ()
13main = do
14 -- Write the methane molecule to a file
15 writeMoleculeToFile "methane.hs" methane
16
17 -- Read the methane molecule from the file
18 molecule <- readMoleculeFromFile "methane.hs"
19
20 -- Print the molecule read from the file
21 print molecule
Listing 14: Haskell code demonstrating the serialisation and writing to disk of a Molecule data type and the reading of a molecule stored in an .hs file, through use of the show and read functions.
An example of the usage is shown.
@@ 6.3 Storage, Transmission, and Readability
This functionality provides a convenient way to save and load molecular data, and is particularly useful when working with large molecules or when molecules are required across different program runs. Although storage and transmission may take place via serialisation into a human and machine-readable string of text, it is important to distinguish between the serialized format of the data and the underlying data type itself. While the ADT can be serialized into a text string, it is crucial to recognize that the ADT itself is not merely a string.
1-- Observed molecule (e.g., H2O)
2observedMolecule :: Molecule
3observedMolecule = h2o
4
5-- A generative model for a molecule
6moleculeModel :: Meas Molecule
7moleculeModel = do
8 let numAtoms = 3
9 atoms <- replicateM numAtoms \$ do
10 symbol <- sample \$ uniformD [C, N, O, H]
11 coord <- sampleCoordinate
12 let attr = elementAttributes symbol
13 let shellsVar = elementShells symbol
14 return \$ Atom { atomID = 0,
15 atomicAttr = attr,
16 coordinate = coord,
17 shells = shellsVar }
18
19 let molecule = Molecule { atoms = atoms, bonds = randomBonds 3 }
20
21 -- Score the model based on the distance to the observed molecule
22 let distance = distMolecule molecule observedMolecule
23 score \$ normalPdf 0 1 distance
24
25 return molecule
26
27-- Sample a coordinate from a normal distribution
28sampleCoordinate :: Meas Coordinate
29sampleCoordinate = do
30 x <- sample \$ normal 0 1
31 y <- sample \$ normal 0 1
32 z <- sample \$ normal 0 1
33 return \$ Coordinate x y z
34
35-- Calculate the Euclidean distance between two coordinates
36euclideanDistance :: Coordinate -> Coordinate -> Double
37euclideanDistance (Coordinate x1 y1 z1) (Coordinate x2 y2 z2) =
38 sqrt \$ (x1 - x2)^2 + (y1 - y2)^2 + (z1 - z2)^2
39
40-- Calculate the distance between two molecules
41distMolecule :: Molecule -> Molecule -> Double
42distMolecule mol1 mol2 =
43 let coords1 = map coordinate (atoms mol1)
44 coords2 = map coordinate (atoms mol2)
45 distances = [euclideanDistance c1 c2 | c1 <- coords1, c2 <- coords2]
46 in minimum distances
47
48-- Example usage
49main :: IO ()
50main = do
51 samples <- mh 0.1 moleculeModel
52 let (molecules, weights) = unzip \$ take 1000 \$ drop 1000 samples
53 putStrLn \$ "Sampled-molecules:" ++ show molecules
54 putStrLn \$ "Weights:" ++ show weights
Listing 15: In a probabilistic programming context, trace Metropolis-Hastings (M-H) operates by sampling entire “traces” of latent variables (i.e., all random choices) from a model [42]. Each step proposes a small perturbation to the current trace, controlled by a jitter parameter, often the standard deviation of the proposal distribution. This parameter must be tuned carefully to ensure efficient exploration of the posterior while maintaining a reasonable acceptance rate. Typically, an initial burn-in period is applied, during which early samples (which have not yet converged) are discarded. This burn-in helps remove the influence of the initial point on the chain, leading to samples that more accurately reflect the stationary distribution of the target posterior. Here the jitter is set to 0.1 and the number of burnin samples is 1000.
6.4 A simple example of probablistic programming
The code block above can be explained as follows:
@@ 7 Discussion
7.1 Limitations
The storage efficiency of the should be further explored. The ADT syntax could result in large file sizes for molecules stored as .hs files, however, the use of list comprehension, and the fact that molfiles and the Molecule data type belong to the same complexity class may mean that the ADT representaion is fact smaller than some mol files, for the same degree of expressiveness. By storing a molecule as function, it is possible to significantly reduce the size of the molecule.
7.2 Future work
The ADT so far presented has the ability to be improved upon, extended, and utilised in empirical experiments in order to verify and test the representation’s utility across a number of domains in cheminformatics.
7.2.1 Augmentation of the ADT
To improve the representational ability of the ADT, it should be augmented to enable representation of tautomerism, ionic bonding, polymerism, and quantum information such as spin, none of which are represented in the current ADT.
Tautomers are in medications from sildenafil (Viagra) [59], to Warfarin, to Remdesivir (used for Covid 19), Tetracyclines and antibiotics. Representation of tautomers could allow exploration of a greater chemical space, and facilitate tasks such as prediction of ligand-protein binding, necessary for many drug design tasks[60, 61].
Polymers are important in many areas of chemistry and materials science. By using lists and recursive data structures, the representation could be augmented to handle polymeric molecules. This extension would enable the modeling and analysis of complex macromolecular systems, opening up new possibilities for polymer design and optimization.
The inclusion of ionic bonding to the ADT would be another significant enhancement. Ionic interactions are critical for the stability and functionality of nucleic acids, proteins, and membranes, as well as for processes like catalysis and ion transport [62, 63, 64]. Augmenting the ADT to represent ions and their bonding explicitly would enable the modeling of charged species and their interactions. This could also facilitate the study of electrochemical properties and enable the design of materials such as ionic liquids and solid electrolytes, which are essential for applications ranging from energy storage to green chemistry [65].
The extension of the ADT to include quantum mechanical properties such as spin and angular momentum represents a significant opportunity. Quantum mechanical features are crucial for accurately modeling the electronic structure and predicting molecular reactivity, as they account for phenomena like electron spin coupling and charge distribution, which play a key role in determining molecular properties [12, 66]. By incorporating these properties into the ADT, it could better capture the behavior of molecules in complex environments, including those encountered in catalysis and photochemistry. Such advancements would also enable the ADT to align more closely with the requirements of quantum chemistry simulations, which often rely on accurate spin and orbital information for predicting reaction mechanisms and designing new materials [11, 1].
7.2.2 Development of powerful, user-friendly library
To make the most of the ADT, a powerful, user-friendly library of functions should to be created. Functions should include the ability to convert between the ADT and other common representaions, which would enable use of the ADT with existing potentially vast chemical databases.
Canonicalisation is also a concern for the ADT, as molecules may have different indicies, and there may be differing atomic coordinates. The permuations of a set of indicies, would however form a group, and the group properties of the ADT can be leveraged for this purpose.
7.2.3 Futher enhancing the ADT
To further strengthen the type safety and invariance guarantees of the representation, exploring dependently typed programming languages, such as Agda, could be highly beneficial. Dependent types allow for the encoding of more sophisticated invariants and constraints directly into the type system, ensuring that only valid and consistent molecular structures are expressible. This increased level of type safety can help catch potential errors at compile-time and provide stronger guarantees about the correctness of the representation.
Further work could explore several avenues to enhance the proposed algebraic data type (ADT) representation and its applications. Integrating Haskell’s Lens library [67] could simplify the manipulation of complex molecular structures by providing a composable, modular and type-safe way to access and modify nested data fields.
To enhance the modularity and composability of the ADT representation, exploring techniques like modular syntax trees and the “Data Types a la Carte” approach [68] could prove highly valuable. By decomposing the molecular representation into smaller, reusable components, we can create a more flexible and adaptable framework. This modular design would allow for the easy integration of new features, the creation of domain-specific languages for molecular manipulation, and the development of reusable libraries and tools for cheminformatics.
By pursuing these avenues of future work, we can further strengthen the capabilities and thus impact the community, encouraging them to use the ADT-based molecular representation, hopefully accelerating the discovery of novel and useful molecules. Implementing backpropagation using differentiable programming [69] would enable using gradient based inference algorithms.
Integrating the ADT with quantum sampling techniques could lead to breakthroughs in predicting molecular properties by efficiently simulating complex quantum systems. However, realizing this goal would require close collaboration between experts in quantum computing, chemical informatics, and molecular modeling to develop efficient and correct quantum algorithms.
7.2.4 Experimental validation of the ADT
At present, the efficiency and efficacy of the ADT in cheminformatics tasks are conjectures. To validate the utility of the ADT as a molecular representation, empirical work is needed to benchmark its performance against representations such as SMILES and SELFIES, on tasks like molecular property prediction (e.g. calculation of logP and logS), virtual screening, and de novo drug design.
Another avenue for future work is utilising the proposed representation with geometric deep learning techniques to exploit symmetries and invariances in molecular structures [2]. By leveraging mathematical properties of ADTs, such as group invariances and equivariances and their verification in Haskell through typeclasses, one could develop neural network architectures that are specifically tailored to the unique characteristics of those molecules. For instance, the choice of index for each atomID used in the representation should not influence a model, given the same molecule but with a different ordering of atomic IDs.
An additional avenue is using equational reasoning to investiage algebraic approaches in molecular modelling [70].
Neural networks can be designed to be invariant to permutations of atomic indexes, ensuring that molecularly equivalent structures are treated identically regardless of labeling of indices or bonds. Similarly, molecular features such as the atomSymbol, are both translation and rotation invariant, while others may only be translation invariant. Incorporating these symmetries directly into the neural network architecture could significantly reduce the amount of training data required and improve the generalization capability of the models, but this would have to be demonstrated.